# Aberdeen Proving Ground, USA

### v2

@@ Abstract
Computationally efficient solutions for pseudometrics quantifying deviation from topological conjugacy between dynamical systems are presented. Deviation from conjugacy is quantified in a Pareto optimal sense that accounts for spectral properties of Koopman operators as well as trajectory geometry. Theoretical justification is provided for computing such pseudometrics in Koopman eigenfunction space rather than observable space. Furthermore, it is shown deriving the pseudometrics from unitary transformations is necessary to recover a value of zero if two systems are topologically conjugate. Therefore the pseudometrics for quantifying deviation from conjugacy are based on analytical solutions for unitary transformations in Koopman eigenfunction space. Finally, geometric considerations for the deviation from conjugac Pareto optimality problem are used to develop scalar pseudometrics that account for all possible solutions given just two Pareto points based. The approach is demonstrated on two example problems; the first being a simple benchmarking problems and the second an engineering example comparing the dynamics of morphological computation of biological nonlinear muscle actuators to simplified ‘man-made’ or bio-inspired approaches. The benefits of considering operator based and trajectory geometry based dissimilarity measures in a unified and consistent formalism were demonstrated. Overall, the deviation from conjugacy pseudometrics provide practical advantages in terms of efficiency and scalability, while maintaining theoretical consistency.
@@
Data-driven dynamics Topological Conjugacy Koopman eigenfunctions
@@ 1 Introduction
Data-driven approaches are increasingly becoming an effective approach for modeling, predicting, controlling, and designing complex nonlinear dynamical systems across a variety of scientific and technical fields brunton_book . Data-driven approaches are especially effective for experimental observations of compex systems for which no accurate equations are available, or analyses for which knowledge of empirical or first-principles governing nonlinear equations must be interrogated in order to understand the system’s behavior. Quantitative comparisons (dissimilarity measures) between nonlinear dynamical systems based on data are particularly important for a variety of machine learning and optimization tasks. It is desirable for data-driven approaches to (a) be computationally efficient such that the approach is scalable to analyzing multiple combinations of systems williams_shape ; lds_distance_opt_algorithm and (b) respect the inherent features of the dynamical system that generated the data bollt_conj ; mezicandbanaszuk:2004 . The first attribute implies an ‘efficient’ quantitative measure where ‘efficient’ is widely accepted to mean it can be computed in polynomial time or better comp_complexity . The second desired attribute requires that the approach be anchored to an appropriate equivalence relationship for nonlinear dynamical systems so that systems that are equivalent return zero dissimilarity when compared. Topological conjugacy wiggins , which provides conditions on both trajectories and operator level attributes, is the equivalence relationship explored here.
@@
There has been extensive work in comparing linear dynamical systems from a both a trajectory geometry and operator perspective. Efficient kernel based approaches were developed in bc_kernel ; kernel_ext . These approaches quantify dissimilarity based on trajectory geometry and follow the behavioral framework perspective to identifying dynamical systems behavior_ds . While the focus was on linear dynamical systems, an extension to nonlinear systems was briefly discussed in bc_kernel . The nonlinear systems formulation from bc_kernel closely resembles a Koopman operator applied_koop perspective. However, system or operator theoretic dissimilarity measures were not considered. A systems oriented perspective to computing ‘distance’ between linear dynamical systems was later taken in 6247929 ; lds_distance_opt_algorithm . It was recognized that restricting alignments to the orthogonal (maximally compact) subgroup of real invertible matrices enables an efficient approach to computing distances. Topological conjugacy was not explored in 6247929 ; lds_distance_opt_algorithm .
@@
Another trajectory focused approach is generalized shape metrics williams_shape , which focused on scalability and computational efficiency. It was also shown how generalized shape metrics relate to (and are often equivalent to) other statistical shape approaches such as canonical correlation analysis cca and representational similarity analysis rsa . Similar to 6247929 ; lds_distance_opt_algorithm , the approach is based on the assumption of orthogonal transformations between data sets. By restricting to orthogonal transformations it was shown in williams_shape that the resulting pseudometrics for comparing systems satisfy subadditivity (i.e. triange inequality) metrics . This was considered an important feature for large scale machine learning applications where inability to satisfy the triangle inequality can lead to inconsitent results. Furthermore, the alignment transformations are solved using the orthogonal one-sided Procrustes solution schonemann which can be computed in polynomial (cubic)time. However, the generalized shape metric approach does not account for the inherent properties of the underlying dynamical system.
@@
Koopman operators were used to develop the data-driven approach dynamic similarity analysis (DSA), which was introduced as an extension to the generalized shape metrics ostrow_DSA . The focus of DSA was to account for system level, rather than trajectory, dissimilarity while maintinaing the advantages of upholding the triangle inequality. As a result, the analysis of williams_shape was adapted to the two-sided Procrustes problem in ostrow_DSA where it was shown that orhtogonal (or special orthogonal) transformations are required to guarantee satisfaction of the triangle inequality. However, because the approach requires solution to a non-convex optimization problem, it is not solvable in polynomial time even if using optimization algorithms described in lds_distance_opt_algorithm . It was shown in ostrow_DSA that DSA recovers the pseudometric from redman2022 ; redman2024 if the data-driven Koopman operators are normal. Otherwise, as in 6247929 ; lds_distance_opt_algorithm , DSA cannot be guaranteed to recover zero dissimilarity for topologically conjugate systems in general due to the assumption of orthogonal transformations in observable space.
@@ 2 Data-Driven Koopman Operator Approximation
For a given discrete time nonlinear dynamical system in state space
@@
where $W$ is a matrix in which each row corresponds to an eigenfunction associated with one of the Koopman eigenvaues. The matrix $W$ is computed by taking the transpose of the left-eigenvectors of $K$ dmd ; edmd . It is assumed here that $W$ is invertible.
@@ 2.1 Observable vector
The Koopman observable functions are chosen by the user. Typically, they consist of functions of physical or application relevance, potentially including system states. Assuming the data is in discretized form, the Koopman observable vector used in this study is $\Psi_{n}=\left[\begin{matrix}1&\Psi_{\mathrm{primary},n}^{T}&\Psi_{\mathrm{aux},n}^{T}\end{matrix}\right]$. The primary observables $\Psi_{\mathrm{primary},n}$ denote quantities of interest to a user, such as system states, control inputs $u$, and nonlinear observable functions $\eta(x,u)$ that are of interest,
@@
Similar to EDMD, the the auxiliary observables are used to accurately approximate the nonlinear behavior in the primary observables. Thus they are a rich set of basis functions for capturing nonlinear behavior. They must be included in identification of the Koopman operator, but may not be of inherent interest to a user. The auxiliary Koopman observables used here are based on the correlation functions given by
@@ 2.2 Identification of the Koopman operator
Given $\Psi_{n+1}$ and $\Psi_{n}$, $K$ is typically identified by the least squares approach dmd ; edmd . Alternativey one could take a Gaussian process machine learning perspective to identify $K$ from the data NIPS2007_66368270 ; GP_ML_book . For example, see masuda2019 ; LIAN2020449 for applications to to identification of Koopman operators. The Gaussian process interpolation approach taken here results in each row of $K$ being a generalized least squares solution to the data, often referred to as a kriging model jon01 . Therefore, the predictors for $\Psi_{n+1}$ as functions of $\Psi_{n}$ are the best linear unbiased predictors jon01 . It was found for the engineering example presented below that the kriging and DMD/EDMD approaches led to similar accuracy but the Gaussian process based Koopman operators had lower norms which can be useful for stability of long term predictions stable_koop . Thus the generalized least squares/kriging interpolation approach was used in this study to identify $K$ from $\Psi_{n+1}$ and $\Psi_{n}$ data. The theory and analytical solutions described next in Sections 3 and 4 apply regardless of one’s preferred approach for identifying $K$.
@@ 3 Topological Conjugacy from the Perspective of Koopman Operators
The conceptual foundation for comparing/classifying systems is that pseudometrics should return zero when two systems are equivalent. For dynamical systems, topological conjugacy serves as the theoretical foundation for equivalence wiggins . If two dynamical systems are topologically conjugate to another, then they are equivalent. Building off the notion of equivalence, one can quantify dissimilarity between dynamical systems by their deviation from conjugacy, i.e. if the deviation from conjugacy between system $a$ and $b$ is less than the deviation between system $b$ and $c$, then between system $a$ and $c$, we can quantitatively say that $a$ is more similar to $b$. A brief description of topological conjugacy and the connection with Koopman based data-driven approaches follows. The deviation from conjugacy pseudometrics are developed in Section 4.
@@
where $\|\;\|_{F}$ is the Frobenius norm. It is assumed that $K_{f}$ and $K_{g}$ have equal dimensionality. Note that translation can be neglected since affine transformations cannot lead to similarity transformations between two linear systems; i.e. $K_{f}\neq h^{-1}K_{g}h$ when $h\circ\Psi_{f}=T\Psi_{f}+b$.
Equations 9 and 10 represent the equivalence conditions that any pseudometric consistent with topological conjugacy should recover if two systems are indeed conjugate. However, before proceeding with the derivation of such pseudometrics, it is first beneficial to recast the conditions in Eqs. 9 and 10 associated with Koopman observable space, $\Psi$-space, into equivalent conditions written in Koopman eigenfunction space, $\Phi$-space. As will be discussed next and in Section 4, the pseudometrics in $\Phi$-space hold several advantages compared to solving for appropriate transformation based on Eqs. 9 and 10.
@@
The equivalence between topological conjugacy conditions in $\Psi$-space and $\Phi$-space established in Proposition 1 provides the conceptual foundation for the development of the pseudometrics in $\Phi$-space. This allows for computing deviations from conjugacy based on $C$ rather than $T$. As will be shown, computationally efficient analytical solutions for $C$ under certain conditions can be found whereas this is not the case if attempting to directly solve for $T$ in $\Psi$-space.
@@ 4 Deviation from Conjugacy Pseudometrics
The development of deviation from conjugacy pseudometrics for comparing dynamical systems is described in this section. Conceptual perspective is provided first, followed by simplifications to the Pareto optimality problem that needs to be solved. Given these simplifications, computationally efficient analytical solutions in $\Phi$-space are provided, followed by their $\Psi$-space representations obtained by appropriate transformation.
@@ 4.1 Deviation from conjugacy
Motivated conceptually by the homeomorphic defect approach bollt_conj , dissimilarity between dynamical systems can be accounted for by quantifying how much the systems deviate from topological conjugacy. The perspective in bollt_conj was based on having a “commuter” $h$ that satisfies Eq. 7 and then quantifying dissimilarity in terms of how $h$ differs from being a homeomoprhism. A somewhat inverse perpsective is taken here in that it is assumed that $h$ is a homeomorphism and dissimalrity is based on the error with respect to Eq. 7. These errors (and their $\Phi$-space instantiations), which are referred to as conjugacy residuals here, serve as vector components in objective function space. Deviations from conjugacy pseudometrics are then defined based on magnitudes of these vectors.
@@
$\begin{split}&\min_{C\in GL(N)}\left(r_{1},r_{2}\right),\\ &r_{1}=\left\|\Phi_{g}-C\Phi_{f}\right\|_{F},\;r_{2}=\left\|\Lambda_{f}-C^{-1}\Lambda_{g}C\right\|_{F}.\;\end{split}$ (18)
@@
The conjugacy residuals can be viewed as vector components in objective functions space, where the origin corresponds to topologically conjugacy. Based on these residuals, deviation from conjugacy is defined as follows:
@@ 4.2.1 Solutions over the group of unitary transformations
The first simplification that is considered is based on the transformation attributes that are required to capture topological conjugacy between two systems. When considering $C\in GL(N_{\Psi})$, this implies transformations that will account for rotation, reflection, and stretching. Thus, an important question is whether all three transformation components are required. This is addressed in Theorem 4.1 below.
@@
where the cyclic prperty of the trace $\operatorname{Tr}$ operation and the trace product idenity $\operatorname{Tr}(AB^{*})=\operatorname{Tr}(BA^{*})$ were used to obtain the second line of Eq. 26.
Proceeding in a similar manner as before, it is assumed without loss of generality that $C=\Gamma P$ where $\Gamma,P\in U(N_{\Psi})$ and $P$ is the same permutation matrix in Eq. 22 that reorders the eigenvaues of system $f$ to match the ordering of eigenvalues in system $g$, i.e. $P\Lambda_{f}P^{\mathrm{T}}=\Lambda_{g}$. Substituting $C=\Gamma P$ into the final term of Eq. 26 leads to
@@
where $\bar{\lambda}_{g,i}$ is the complex conjugate of the $i$th eigenvalue of system $g$ and $\bar{\Gamma}_{ii}$ is the complex conjugate of element $(i,i)$ of $\Gamma$. The cyclic property of $\operatorname{Tr}$ and that fact that $P\Lambda_{f}^{*}P^{\mathrm{T}}=\Lambda_{g}^{*}$ were used to obtain the second line of Eq.27, and Eq. 8 was used to obtain the final line. From the last line of Eq. 27 it is clear that topological conjugacy occurs when $\Gamma_{ii}\bar{\Gamma}_{ii}=1$. Since $\Gamma$ is unitary, this condition implies $\Gamma$ is a any diagonal matrix where each diagonal element has magnitude equal to one. Therefore, the proof of Theorem 4.1 is completed by choosing $\Gamma=\Theta$ such that the same unitary matrix $C$ satisfies both $\left\|\Lambda_{f}-C^{*}\Lambda_{g}C\right\|_{F}=0$ and $\left\|\Phi_{g}-C\Phi_{f}\right\|_{F}=0$.
@@
$\begin{split}&\min_{C\in U(N)}\left(r_{1},r_{2}\right),\\ &r_{1}=\left\|\Phi_{g}-C\Phi_{f}\right\|_{F},\;r_{2}=\left\|\Lambda_{f}-C^{*}\Lambda_{g}C\right\|_{F}.\end{split}$ (28)
Theorem 4.1 implies that Eq. 28 can be expected to be accurate when two systems are close to conjugate since stretching is an unnecessary transformation component at conjugacy. Indeed, stretching would be expected to be play a more significant role in precise quantification of deviation when two systems are far from topological conjugacy. However, in such cases, Eq. 28 would register large values of $r_{1}$ and/or $r_{2}$ thus capturing the relative argument that the two systems are very different from the perspective of topological conjugacy.
@@
When solving Eq. 9 and Eq. 10 in $\Psi$-space by restricting $T$ to be unitary(or orthogonal), e.g. 6247929 ; lds_distance_opt_algorithm ; williams_shape ; ostrow_DSA , the resulting pseudometrics will not yield zero dissimilarity for topologically conjugate systems in general. Since conjugacy residuals satisfy Theorem 4.1, the pseudometrics in this study will recover zero when $f$ and $g$ are topologically conjugate. Therefore the computational efficiency of associated with the approaches in 6247929 ; lds_distance_opt_algorithm ; williams_shape ; ostrow_DSA are achieved while maintaining theoretical consitency with topological conjugacy by computing residuals in $\Phi$-space.

### v4

@@ Abstract
Computationally efficient solutions for pseudometrics quantifying deviation from topological conjugacy between dynamical systems are presented. Deviation from conjugacy is quantified in a Pareto optimal sense that accounts for spectral properties of Koopman operators as well as trajectory geometry. Theoretical justification is provided for computing such pseudometrics in Koopman eigenfunction space rather than observable space. Furthermore, it is shown that theoretical consistency with topological conjugacy can be maintained when restricting the search for optimal transformations between systems to the unitary group. Therefore the pseudometrics are based on analytical solutions for unitary transformations in Koopman eigenfunction space. Geometric considerations for the deviation from conjugacy Pareto optimality problem are used to develop scalar pseudometrics that account for all possible optimal solutions given just two Pareto points. The approach is demonstrated on two example problems; the first being a simple benchmarking problem and the second an engineering example comparing the dynamics of morphological computation of biological nonlinear muscle actuators to simplified mad-made (including bioinspired) approaches. The benefits of considering operator and trajectory geometry based dissimilarity measures in a unified and consistent formalism is demonstrated. Overall, the deviation from conjugacy pseudometrics provide practical advantages in terms of efficiency and scalability, while maintaining theoretical consistency.
@@
Data-driven dynamics Topological conjugacy Dissimilarity measures Koopman eigenfunctions
@@ 1 Introduction
Data-driven approaches are becoming increasingly prevalent for modeling, predicting, controlling, and designing complex nonlinear dynamical systems across a variety of scientific and technical fields brunton_book . Data-driven approaches are especially effective for analyzing experimental observations of complex dynamical systems for which no accurate equations are available, or analyses for which knowledge of empirical or first-principles governing nonlinear equations must be interrogated in order to understand the system’s behavior. Quantitative comparisons (dissimilarity measures) between nonlinear dynamical systems based on data are particularly important for a variety of machine learning and optimization tasks. As a result, it is desirable for data-driven dissimilarity measures to (a) be computationally efficient such that the approach is scalable to analyzing multiple combinations of systems williams_shape ; lds_distance_opt_algorithm and (b) respect the inherent features of the dynamical system that generated the data bollt_conj . The first attribute implies an ‘efficient’ quantitative measure where ‘efficient’ is widely accepted to mean it can be computed in polynomial time or better comp_complexity . The second desired attribute requires that the approach be anchored to an appropriate equivalence relationship for nonlinear dynamical systems so that systems that are equivalent return zero dissimilarity when compared. Topological conjugacy wiggins , which provides conditions on both trajectories and operator level dissimilarity, is the equivalence relationship explored here.
@@
There has been extensive work in comparing linear dynamical systems from both a trajectory geometry and operator perspective. Efficient kernel based approaches were developed in bc_kernel ; kernel_ext . These approaches quantify dissimilarity based on trajectory geometry and follow the behavioral framework perspective to identifying dynamical systems behavior_ds . While the focus was on linear dynamical systems, an extension to nonlinear systems was briefly discussed in bc_kernel . The nonlinear systems formulation from bc_kernel closely resembles a Koopman operator applied_koop perspective. However, system or operator theoretic dissimilarity measures were not considered. A systems oriented perspective to computing ‘distance’ between linear dynamical systems was later taken in 6247929 ; lds_distance_opt_algorithm . It was recognized that restricting to the orthogonal (maximally compact) subgroup of real invertible matrices could reduce the computational expense of solving for optimal alignments between systems. Topological conjugacy considerations were not explored in 6247929 ; lds_distance_opt_algorithm . Furthermore, generalizations to complex transformations and the relevance to dynamical systems was not studied.
@@
Another trajectory focused approach is generalized shape metrics williams_shape , which focused on scalability and computational efficiency. It was also shown how generalized shape metrics relate to (and are often equivalent to) other statistical shape approaches such as canonical correlation analysis cca and representational similarity analysis rsa . Similar to 6247929 ; lds_distance_opt_algorithm , the approach is based on the assumption of orthogonal transformations between data sets. By restricting to orthogonal transformations it was shown in williams_shape that the resulting pseudometrics for comparing systems satisfy subadditivity (i.e. triange inequality) metrics . This was considered an important feature for large scale machine learning applications where inability to satisfy the triangle inequality can lead to inconsitent results. Furthermore, the alignment transformations are solved using the orthogonal one-sided Procrustes solution schonemann which can be computed in polynomial (cubic) time. However, the generalized shape metric approach does not account for the inherent properties of the underlying dynamical system operator.
@@
Koopman operators served as the foundation for development of dynamic similarity analysis (DSA)ostrow_DSA , which was introduced as an extension to generalized shape metrics. The focus of DSA was on accounting for system level, rather than trajectory level, dissimilarity while maintaining the advantages of upholding the triangle inequality. As a result, the analysis of williams_shape was adapted to the two-sided Procrustes problem in ostrow_DSA where it was shown that orthogonal (or special orthogonal) transformations guarantee satisfaction of the triangle inequality. However, because the approach requires solution to a non-convex optimization problem, it is not solvable in polynomial time even if using fast optimization algorithms such as the one described in lds_distance_opt_algorithm . It was shown in ostrow_DSA that DSA recovers the pseudometric from redman2022 ; redman2024 if the data-driven Koopman operators are normal. Otherwise, as in 6247929 ; lds_distance_opt_algorithm , DSA cannot be guaranteed to recover zero dissimilarity for topologically conjugate systems in general due to the assumption of orthogonal transformations in observable space.
@@ 2 Data-Driven Koopman Operator Approximation
For a nonlinear dynamical system in which data is obtained in discrete time, one has
@@
where $W$ is a matrix in which each row corresponds to an eigenfunction associated with one of the Koopman eigenvaues. The matrix $W$ is computed by taking the transpose of the left-eigenvectors of $K$ dmd ; edmd . It is assumed here that $W$ is invertible and $K$ is diagonalizable.
@@ 2.1 Observable vector
The Koopman observable functions are chosen by the user. Typically, they consist of functions of physical or application relevance, potentially including system states. Assuming the data is in discretized form, the Koopman observable vector used in this study is $\Psi_{n}=\left[\begin{matrix}1&\Psi_{\mathrm{primary},n}^{T}&\Psi_{\mathrm{aux},n}^{T}\end{matrix}\right]$. The primary observables $\Psi_{\mathrm{primary},n}$ denoting quantities of interest are chosen here to be the system states, control inputs $u$, and certain nonlinear observable functions $\eta(x,u)$,
@@
Similar to EDMD, the auxiliary observables are used to accurately approximate the nonlinear behavior in the primary observables. Thus they are a rich set of basis functions for capturing nonlinear behavior. They must be included in identification of the Koopman operator, but may not be of inherent interest to a user. The auxiliary Koopman observables used here are based on the correlation functions given by
@@ 2.2 Identification of the Koopman operator
Given $\Psi_{n+1}$ and $\Psi_{n}$, $K$ is typically identified by the least squares approach dmd ; edmd . Alternativey one could take a Gaussian process machine learning perspective to identify $K$ from the data NIPS2007_66368270 ; GP_ML_book . For example, see masuda2019 ; LIAN2020449 for applications to to identification of Koopman operators. The Gaussian process interpolation approach taken here results in each row of $K$ being a generalized least squares solution to the data, often referred to as a kriging model jon01 . Therefore, the predictors for $\Psi_{n+1}$ as functions of $\Psi_{n}$ are the best linear unbiased predictors jon01 . It was found for the engineering example presented in Section 5 that the kriging and DMD/EDMD approaches led to similar accuracy but the Gaussian process based Koopman operators had lower norms which can be useful for stability of long term predictions stable_koop . Thus the generalized least squares/kriging interpolation approach was used in this study to identify $K$ from $\Psi_{n+1}$ and $\Psi_{n}$ data. The theory and analytical solutions described next in Sections 3 and 4 apply regardless of one’s preferred approach for identifying $K$.
@@ 3 Topological Conjugacy from the Perspective of Koopman Operators
The conceptual foundation for comparing/classifying systems is that pseudometrics should return zero when two systems are equivalent. For dynamical systems, topological conjugacy serves as the theoretical foundation for equivalence wiggins . If two dynamical systems are topologically conjugate to another, then they are equivalent. Building off the notion of equivalence, one can quantify dissimilarity between dynamical systems by their deviation from conjugacy, i.e. if the deviation from conjugacy between system $a$ and $b$ is less than the deviation between system $b$ and $c$, then among systems $a$ and $c$, we can quantitatively say that $a$ is more similar to $b$. A brief description of topological conjugacy and the connection with Koopman based data-driven approaches follows. The deviation from conjugacy pseudometrics are developed in Section 4.
@@
where $\|\;\|_{F}$ is the Frobenius norm. Note that translation can be neglected since affine transformations cannot lead to similarity transformations between two linear systems; i.e. $K_{f}\neq h^{-1}K_{g}h$ when $h\circ\Psi_{f}=T\Psi_{f}+b$.
Equations 9 and 10 represent the equivalence conditions that any pseudometric consistent with topological conjugacy should recover if two systems are indeed conjugate. However, before proceeding with the derivation of such pseudometrics, it is first beneficial to recast the conditions in Eqs. 9 and 10 associated with Koopman observable space, $\Psi$-space, into equivalent conditions written in Koopman eigenfunction space, $\Phi$-space. As will be discussed next and in Section 4, the pseudometrics in $\Phi$-space hold several advantages compared to solving for optimal transformations based on Eqs. 9 and 10.
@@
The equivalence between topological conjugacy conditions in $\Psi$-space and $\Phi$-space established in Proposition 1 provides the conceptual foundation for the development of the pseudometrics in $\Phi$-space. This allows for computing deviations from conjugacy based on $C$ rather than $T$. As will be shown, computationally efficient and theoretically consistent analytical solutions for $C$ under certain group restrictions can be found whereas this is not the case if attempting to directly solve for $T$ in $\Psi$-space.
@@ 4 Deviation from Conjugacy Pseudometrics
The development of deviation from conjugacy pseudometrics for comparing dynamical systems is described in this section. Conceptual perspective is provided first, followed by simplifications to the Pareto optimality problem that needs to be solved. Given these simplifications, computationally efficient analytical solutions in $\Phi$-space are provided, followed by their corresponding $\Psi$-space representations.
@@ 4.1 Deviation from conjugacy
Motivated conceptually by the homeomorphic defect approach bollt_conj , dissimilarity between dynamical systems can be accounted for by quantifying how much the systems deviate from topological conjugacy. The perspective in bollt_conj was based on having a “commuter” $h$ that satisfies Eq. 7 and then quantifying dissimilarity in terms of how $h$ differs from being a homeomoprhism. A somewhat inverse perpsective is taken here in that it is assumed that $h$ is a homeomorphism and dissimalrity is based on the error with respect to Eq. 7. These errors (and their $\Phi$-space instantiations), which are referred to as conjugacy residuals here, serve as vector components in an objective function space. Deviations from conjugacy pseudometrics are then defined based on magnitudes of these vectors.
@@
$\begin{split}&\min_{C\in GL(N_{\Psi})}\left(r_{1},r_{2}\right),\\ &r_{1}=\left\|\Phi_{g}-C\Phi_{f}\right\|_{F},\;r_{2}=\left\|\Lambda_{f}-C^{-1}\Lambda_{g}C\right\|_{F}.\;\end{split}$ (18)
@@
The conjugacy residuals can be viewed as vector components in objective function space, where the origin corresponds to topologically conjugacy. Based on these residuals, deviation from conjugacy is defined as follows:
@@ 4.2.1 Solutions over the group of unitary transformations
The first simplification that is considered is based on the transformation attributes that are required to capture topological conjugacy between two systems. When considering $C\in GL(N_{\Psi})$, this implies transformations that will account for rotation, reflection, and stretching. Thus, an important question is whether all three transformation components are necessary? This is addressed in Theorem 4.1 below.
@@
where the cyclic prperty of the trace $\operatorname{Tr}$ operation and the trace product idenity $\operatorname{Tr}(AB^{*})=\operatorname{Tr}(BA^{*})$ were used to obtain the second line of Eq. 26. Proceeding in a similar manner as before, it is assumed without loss of generality that $C=\Gamma P$ where $\Gamma,P\in U(N_{\Psi})$ and $P$ is the same permutation matrix in Eq. 22 that reorders the eigenvaues of system $f$ to match the ordering of eigenvalues in system $g$, i.e. $P\Lambda_{f}P^{\mathrm{T}}=\Lambda_{g}$. Substituting $C=\Gamma P$ into the final term of Eq. 26 leads to
@@
where $\bar{\lambda}_{g,i}$ is the complex conjugate of the $i$th eigenvalue of system $g$ and $\bar{\Gamma}_{ii}$ is the complex conjugate of element $(i,i)$ of $\Gamma$. The cyclic property of $\operatorname{Tr}$ and that fact that $P\Lambda_{f}^{*}P^{\mathrm{T}}=\Lambda_{g}^{*}$ were used to obtain the second line of Eq. 27, and Eq. 8 was used to obtain the final line. From the last line of Eq. 27 it is clear that topological conjugacy occurs when $\Gamma_{ii}\bar{\Gamma}_{ii}=1$. Since $\Gamma$ is unitary, this condition implies $\Gamma$ is any diagonal matrix where each diagonal element has magnitude equal to one. Therefore, the proof of Theorem 4.1 is completed by choosing $\Gamma=\Theta$ such that the same unitary matrix $C$ satisfies both $\left\|\Lambda_{f}-C^{*}\Lambda_{g}C\right\|_{F}=0$ and $\left\|\Phi_{g}-C\Phi_{f}\right\|_{F}=0$.
@@
$\begin{split}&\min_{C\in U(N_{\Psi})}\left(r_{1},r_{2}\right),\\ &r_{1}=\left\|\Phi_{g}-C\Phi_{f}\right\|_{F},\;r_{2}=\left\|\Lambda_{f}-C^{*}\Lambda_{g}C\right\|_{F}.\end{split}$ (28)
Theorem 4.1 implies that Eq. 28 can be expected to be quantitatively accurate when two systems are close to conjugate since stretching is an unnecessary transformation component at conjugacy. Indeed, stretching would be expected to be play a more significant role in precise quantification of deviation when two systems are very far from topological conjugacy. However, in such cases, Eq. 28 would register large values of $r_{1}$ and/or $r_{2}$ thus capturing the relative argument that the two systems are very different from the perspective of topological conjugacy. Thus Eq. 28 is expected to be qualitatively accurate very far from conjugacy.
@@
When solving Eq. 9 and Eq. 10 in $\Psi$-space by restricting $T$ to be unitary(or orthogonal), e.g. 6247929 ; lds_distance_opt_algorithm ; williams_shape ; ostrow_DSA , the resulting pseudometrics will not yield zero dissimilarity for topologically conjugate systems in general. Since conjugacy residuals satisfy Theorem 4.1, the pseudometrics in this study will recover zero when $f$ and $g$ are topologically conjugate. Therefore the computational efficiency associated with the approaches in 6247929 ; lds_distance_opt_algorithm ; williams_shape ; ostrow_DSA are achieved while maintaining theoretical consistency with topological conjugacy by computing residuals in $\Phi$-space.